### v2

# SEFL: Enhancing Educational Assignment Feedback with LLM Agents

Abstract
Providing high-quality feedback to student assignments is crucial for student success, but it is constrained by time and costs. In this work, we introduce Synthetic Educational Feedback Loops (SEFL), a synthetic data framework designed to generate data that resembles immediate, on-demand feedback at scale without relying on extensive, real-world student assignments. To get this type of data, two large language models (LLMs) operate in teacher–student roles to simulate assignment completion and formative feedback, generating synthetic pairs of student work and corresponding critiques and actionable improvements from a teacher. With this data, we fine-tune smaller, more computationally efficient LLMs on these synthetic pairs, enabling them to replicate key features of high-quality, goal-oriented feedback. Unlike personalized tutoring approaches that offer multi-turn, individualized instruction, SEFL specifically focuses on replicating the teacherstudent assignment feedback loop in higher education. Through comprehensive evaluations with four LLM judges and three human experts, we demonstrate that SEFL-tuned models outperform both their non-tuned counterparts in feedback quality and an existing baseline. The potential for societal impact is reinforced by extensive qualitative comments by ratings by human stakeholders—both students and higher education instructors. All in all, SEFL has substantial potential to transform feedback processes for higher education and beyond.
Code — https://github.com/jjzha/sefl
Datasets and Models — https://huggingface.co/collections/jjzha/sefl-synthetic-educational-feedback-loops-67b48768dab5123a2e3e9d69
1 Introduction
Constructive feedback is a cornerstone of higher education, promoting critical thinking and fostering deeper understanding (Hattie 2008; Costello and Crane 2013). In higher education settings, however, providing consistent, high-quality feedback is complicated by privacy, consent, and transparency considerations in data collection (Fischer et al. 2020; Suresh et al. 2022; Demszky and Hill 2023; Wang and Demszky 2024; Wang et al. 2024b; Lindsay et al. 2024), in addition to being a labor-intensive task. Advances in language technology offer opportunities to automate and augment feedback processes, addressing these limitations.
In particular, LLMs have shown progress in education (Wang et al. 2024c), including automated grading (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024) and personalized tutoring (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024, 2025; Wang et al. 2024a). Yet, automating teacher–student assignment feedback with LLMs remains an open question. We seek to answer: RQ. How can synthetic teacher–student interactions generated by LLMs be leveraged to enable scalable and effective educational student assessment feedback?
Here, we introduce Synthetic Educational Feedback Loops (SEFL), a framework that generates synthetic teacher–student interactions using LLM agents. In this framework, two LLMs, one acting as the teacher and the other as the student, simulating formative feedback workflows (Conole and Oliver 2006; Nicol 2007). To induce better feedback from LLMs this synthetic data from the agents is then used to fine-tune smaller autoregressive language models, allowing the development of scalable educational feedback systems that can operate efficiently on modest computational infrastructure, available in higher education institutions without the need for privacy sensitive data.
We present several findings. (i) Through empirical and qualitative analysis of expert annotator comments, we find that larger models tend to provide more actionable, goal-oriented, and user-friendly feedback. They are also more consistent and better at supporting student autonomy, showing the potential for use in real-world contexts. (ii) Empirically, SEFL-tuned models outperform their non-tuned versions in win rate evaluations by 4 LLM judges and 3 human experts in giving assessment feedback. (iii) We observe strong agreement among human evaluators regarding feedback quality. (iv) We compare our approach to Book2Dial (Wang et al. 2024a), showing that SEFL achieves a higher win rate according to 3 of the 4 LLM judges.
Contributions.
We contribute the following in this work:
SEFL: A framework that simulates teacher student feedback loops with paired language model agents.
19,841 assignment–feedback pairs generated by SEFL to fine-tune smaller language models.
A comprehensive mix of human and language model evaluation that highlights strengths and limitations of SEFL with extensive qualitative analysis of the feedback provided by human experts.
An open source release of all models, code, and data.
2 Related Work
NLP & Education.
Large language models are now supporting a broad spectrum of educational tasks. In automated grading, they score short answers, essays, and even programming assignments with accuracy that approaches expert instructors, which eases faculty workload and releases time for mentoring (Ke and Ng 2019; Ramesh and Sanampudi 2022; Stahl et al. 2024). For personalized tutoring, conversational agents powered by these models adapt explanations, hints, and examples to each learner’s background knowledge and preferred style, producing measurable gains in engagement and achievement (Yun et al. 2024; Liu et al. 2024c; Rooein and Hovy 2024; Ross and Andreas 2024; Kwon et al. 2024; Zhang et al. 2024). Research on peer learning shows that the same technology can mediate small-group discussions, suggest prompts, and highlight diverse viewpoints, leading to richer collaboration (Bauer et al. 2023). In mathematics, aligning word problems and proofs to grade-level objectives has been automated with encouraging results, reducing the manual effort needed to curate question banks (Botelho et al. 2023). Critical thinking curricula benefit as well; LLMs can challenge students to justify claims, detect fallacies, and refine arguments in real time (Guerraoui et al. 2023). The models have also begun to assist scholars: studies report successful use for screening literature, summarizing drafts, and aligning reviewer comments with revision plans (Liang et al. 2024; Sonkar et al. 2024). Complementing these functions are analytics tools that track learning trajectories and surface early warnings when a student slips behind (Schwarz et al. 2018; Aslan et al. 2019; Alrajhi et al. 2021).
Despite this growing body of work, prior studies have rarely targeted the systematic generation of rich feedback on open-ended student submissions and assignments at scale. The present study addresses that gap by using LLMs to create extensive comment sets that teachers can accept as-is or trim to taste. Decades of scholarship define effective feedback as goal-oriented, actionable, timely, user-friendly, and consistent while fostering self-evaluation (Carless et al. 2011; Wiggins 2012). Feedback that rambles can confuse learners, so concise wording is usually preferable, and comments that arrive soon after the original effort drive steady improvement (Wiggins 2012). By producing immediate responses and tailoring suggestions to rubric criteria, LLM based systems stand well positioned to satisfy these guidelines while operating at classroom and institution scale.
Synthetic Data Frameworks.
Recent research shows how collaborative agentic LLMs can synthesize large-scale interactional datasets for educational tasks. For example, CAMEL (Li et al. 2023) uses cooperative role-based dialogues to achieve shared objectives, while SimSeek (Kim et al. 2022) uses agent-based conversations to build comprehensive information-seeking datasets. In education, SocraticLM (Liu et al. 2024b) simulates Socratic tutoring through multi-turn dialogue, and Book2Dial (Wang et al. 2024a) generates teacher-student conversations from textbooks. In contrast, SEFL focuses on concise teacher-student feedback loops rather than extended instructional dialogues. While Nair et al. (2024) explore iterative revisions, SEFL generates diverse feedback pairs from assignment-answer-feedback tuples, enabling fine-tuning of smaller, cost-effective models for large-scale use.
3 Synthetic Educational Feedback Loops
3.1 Synthetic Data Generation
We use a two-agent framework (Wu et al. 2023) to simulate a feedback loop as in higher education. Both the teacher and student roles are simulated by two separate Llama-3.1-70B models for a two-turn conversation.111Note that if we mention a model, it is always the post-trained version (i.e., -Instruct). The models are tasked to generate assignmentanswerfeedback tuples. First, the student-agent asks for an assignment using Fineweb-Edu (Lozhkov et al. 2024) texts (Figure 1), known for its educational content based on LLM judgments. Second, the teacher-agent creates an assignment that can be of any domain, e.g., math, humanities, role-playing. Then, the student-agent ( ) submits assignments containing a number of explicit errors, and the teacher-agent ( ) provides feedback. We investigate both Qwen2.5-72B and Llama-3.1-70B for interactions. First, we generate 5,000 interaction tuples with each model, where we validate the output as an initial step to investigate initial feedback quality.
We show in Table 1 the results of this first experiment. Out of 5,000 generated examples, Llama-3.1-70B generates 2,513 valid examples (i.e., valid JSON format and each feedback point refers to an error) compared to Qwen2.5-72B with 454 valid examples. For a further check, we use BERTScore (Zhang et al. 2020) as a proxy to see whether each error-feedback pair of the valid generations relate to each other.222We only calculate it of the samples where both error and feedback have the same number of generations. We show that, regardless of Llama-3.1-70B generating more valid examples, the BERTScore (0.877) stays in a similar range as Qwen2.5-72B (0.919); in both cases indicating high similarity. As a last qualitative check, we experimented with several prompts and consolidated into a final prompt found in the supplementary material. Finally, we use Llama-3.1-70B-generated data as the basis for all subsequent model fine-tuning as it generated more valid examples.
| Valid (/ 5,000) | BERTScore | |
| Llama-3.1-70B | 2,513 | 0.877 |
| Qwen2.5-72B | 454 | 0.919 |
Statistics.
After the 5K generated examples, we continue generating example pairs and end up with 19.8K teacher–student feedback pairs. In Table 2, we present the final dataset statistics. We highlight that the generation lengths for each agent are intentionally kept concise (170 subword tokens), based on the hypothesis that overly lengthy feedback may be counterproductive, especially for short assignments. This is in line with observations from Ferguson (2011), who finds that students tend to favor brief comments. We argue that balancing supportive and critical feedback is crucial as, by default, LLMs often produce excessively verbose responses, which can influence the preferences of both humans and language models (Saito et al. 2023).
4 Methodology
4.1 Fine-Tuning Large Language Models
We divide the data into 17,856 training examples and 1,985 validation examples. To test the generalizability of our approach across model scales, we fine-tune five open-weight models, namely Qwen2.5-0.5B, Llama-3.2-1B, Llama-3.2-3B, Llama-3.1-8B, and Qwen2.5-14B on this corpus. The compute we train the models on are AMD Radeon Instinct MI250X GPUs and it took a total of 467 GPU hours.
Supervised Objective.
To fine-tune the models, formally, for each prompt and target sequence we minimize the token-level cross-entropy
where masks out the prompt tokens and activates the loss on reference tokens only.
Optimization Details.
All models train for three epochs with a global batch size of 16 and context lengths of 131K for Qwen2.5 and 128K for the Llama variants. We use AdamW with , (Loshchilov and Hutter 2019), , weight decay , and gradient clipping at norm . The learning rate peaks at after a linear warm-up covering the first five percent of steps and then follows a linear decay.
| Feature | Value |
| Instances | 19,841 |
| Assignment Length | 78.6 |
| Length (Student Agent) | 168.1 |
| # Errors Points | 2.5 |
| Length # Errors | 20.7 |
| Length (Teacher Agent) | 120.5 |
| # Feedback Points | 2.5 |
| Length # Feedback | 34.6 |
4.2 Multi-faceted Evaluation
Human Evaluation.
After fine-tuning the language models with the assignment–feedback pairs, to test the performance of SEFL, we do human evaluation. We randomly sample 150 instances from the validation set. For each item, both the original instruction tuned model (A) and the model further fine-tuned with SEFL (B) produce feedback. Three human experts compared pairs of feedback responses produced for the same assignment and answer. For each item they read the original prompt, the student submission, and the two candidate feedback drafts from the non-tuned and SEFL-tuned model. Then, they select the feedback, from model A or B, that is better based on four base criteria:
Accuracy. The generated feedback text refers to concrete strengths and weaknesses in the student work and avoid superficial remarks.
Actionability. Suggestions are clear, specific, and realistic for a student to apply.
Conciseness. Wording is brief and focused, with little repetition.
Tone. Language stays constructive and professional while recognizing good elements.
Raters were reminded to value efficiency over length, to prefer targeted advice over general principles, and to ignore formatting tricks. They recorded their choice as A or B and could leave an optional free-text comment. With this we calculate win rate (i.e., the percentage of choosing one feedback text over the other). This has become a de facto standard to evaluate long form text against each other (e.g., Rafailov et al. 2023).
Each row took at most ten minutes, and the guidelines stressed taking regular breaks to sustain attention. We deliberately remove the tie option because a forced choice gives more informative labels and reduces hesitation, while a separate checkbox still lets raters mark assignmentanswerfeedback tuples as unrelated or nonsensical. This happened around 12% of the time, especially in the smaller, less capable models. The full annotation guidelines reported in the supplementary material in the code repository.
Our human raters are in the aged 20–40 and from Europe. One identifies as female and the other two identifies as male. One female and male have a background in Computer Science and one male in Education Engineering. All have extensive experience in teaching and supervision or being taught and supervised, they all work in higher education (at different levels, i.e., research assistant and assistant professors) with near-native English proficiency.
LLM-as-a-Judge.
We also evaluate the fine-tuned models’ output using a LLM-as-a-judge framework, a method gaining traction as a method for evaluating free text output (Liu et al. 2023; Zheng et al. 2024; Chen et al. 2023; Verga et al. 2024; Törnberg 2023; Naismith, Mulcaire, and Burstein 2023; Gilardi, Alizadeh, and Kubli 2023; Kocmi and Federmann 2023; Huang et al. 2024; Gu et al. 2024; Falk et al. 2025). The same 150 random instances are rated by the four LLMs, namely GPT-4o (Hurst et al. 2024), Claude3.5-Sonnet, Command-R+, and DeepSeek-V3 (Liu et al. 2024a). For the closed-source models’ LLM-as-a-judge experiments, we use their respective APIs and the total costs were approximately 10 USD.
For every example the judge model receives the assignment prompt together with the two candidate feedback drafts. It is asked to decide which draft is better on the same four base criteria as the humans. The instruction forbids numeric grades or explanations and requires the judge to output exactly one character, A or B, producing a clean pairwise preference label. We report the full prompt in the supplementary material in the code repository.
5 Results
In Figure 2, we show the win rates of models fine-tuned with SEFL vs. their non-tuned version, evaluated by both humans and LLM-based judges. A value above 50% indicates that the SEFL-tuned models are preferred over their original versions. We show an example of the feedback in Figure 4, where we depict the abridged prompt and feedback by two tuned an non-tuned models.
Human Assessment.
Overall, human rater evaluations in Figure 2 (H#) show that the SEFL-tuned models often attain high win rates, surpassing 90% in several cases with respect to the smaller models. The human annotators differed in their views on the 8B model’s output quality; however, they generally converged on the observation that the fine-tuned 14B model produces superior feedback compared to its original version. By contrast, models not fine-tuned with SEFL had lower win rates, suggesting that SEFL provides an edge in generating more coherent and context-relevant feedback. In addition, we asked annotators whether the synthetic assignmentanswerfeedback sequences were consistent. In over 75% of cases, they confirmed the alignment between assignment, student response, and the feedback given, showing positive contextual relevance.
LLM-as-a-Judge Results.
For the LLM-as-a-judge evaluations (J#) in the same figure, we observe notable differences in win rates depending on the model and scale. The results largely mirror the human assessment trend up to the 3B scale. The results from the four LLM judges (J1: GPT-4o, J2: Claude-3.5-Sonnet, J3: Command-R+, J4: Deepseek-v3) reveal that SEFL-tuned models show varying levels of performance relative to their vanilla counterparts.333Models are picked based on their recency and performance on RewardBench (Lambert et al. 2024), JudgeBench (Tan et al. 2024), and JudgeArena (AtlaAI 2025). For instance, Qwen2.5-0.5B achieved the highest win rates across all four judges (62% on J3), indicating a consistent preference for the fine-tuned version. In contrast, larger models such as Llama-3.1-8B and Qwen2.5-14B exhibit lower win rates, particularly on J3 (16% and 10%, respectively), suggesting that fine-tuning with SEFL may yield diminishing returns or challenges at larger scales. The full judge prompt can be found in the supplementary material.
Agreement.
In Figure 3, we present the pairwise Cohen’s values (Cohen 1960) computed between each language-model judge and the human raters, to observe whether humans and LLM judges agree on which model gives better feedback. The agreement among humans was moderate to substantial: H1 and H3 reached , H1 and H2 , and H2 and H3 (Landis and Koch 1977). Among the models, Claude aligns most closely with both the other judges and the humans; DeepSeek follows, and GPT-4o shows the weakest match. Across all model and human pairs the coefficients range from to , which underlines the subjectivity of feedback evaluation. Command R+ sits at the lower extreme, returning values between and when compared with human raters and with other judges, which signals virtually no agreement. This signals that the human experts mainly agree on the quality of feedback, but for LLM judges this can differ. Overall, there are improvements to be made in terms of agreement between LLMs and humans in terms of feedback quality.
[…]
| 0.5B | “The feedback repeats the assignment rather than pointing out issues.” | “A is concise but misses important detail.” | “Comments do not reflect missing parts of the answer.” |
| 0.5B-SEFL | “Feedback spots deliberate errors and offers clear fixes.” | “A is concise and clear, suggestions align with the rubric.” | “Advice notes that the discuss part is missing and proposes adding it.” |
| 1B | “Feedback echoes what the answer already states.” | “More precise.” | “Comments ignore omitted sections.” |
| 1B-SEFL | “The note opens with strengths and gives concrete next steps.” | “Hints link directly to rubric points.” | “Advice adds two practical examples to guide revision.” |
| 3B | “Both models are good, A sounds nicer.” | “Both are good here.” | “Feedback covers content but wording feels harsh.” |
| 3B-SEFL | “Feedback keeps an encouraging tone while giving actionable points.” | “Structure and tone feel balanced.” | “Comments cover content, style, and structure in one cohesive note.” |
| 7B | “Feedback addresses surface errors but misses deeper reasoning.” | “Detailed but drifts off the prompt.” | “Overlooks the rubric item about evidence.” |
| 7B-SEFL | “Points out reasoning gaps and proposes specific fixes.” | “Stays focused on the prompt and remains concise.” | “Directly cites the missing evidence section.” |
| 14B | “Feedback is thorough but wording is opaque.” | “Long sentences reduce clarity.” | “Some comments repeat earlier points.” |
| 14B-SEFL | “Clear and student friendly while retaining depth.” | “Uses short sentences and connects advice to rubric.” | “Groups comments by theme which avoids repetition.” |
| Aspect | Base Model | SEFL-tuned Model | Preferred |
| Consistency | “Feedback from model A fails to address key aspects of the answer, such as suddenly changing the name of the main character.” | “Not an answer but A properly identified it!” | SEFL |
| Goal orientation | “In many cases, answers are shorter than required. Not reflected in feedback.” | “Finally, B finds that the answer is incomplete.” | SEFL |
| Tone | “Model B: Harsh tone.” | “Both models are good, but model A is nicer in tone and actionability.” | SEFL |
| User friendliness | “Model B is best, but is way too elaborate.” | “B is more clear and concise.” | SEFL |
| Actionability | “Model A also provides partial solution.” | “Model B is not accurate and provides an answer instead of feedback.” | Base |
6 Discussion
6.1 Human Qualitative Insights
In addition to the win rates in Figure 2, our human annotators provided rich qualitative feedback on the model outputs, which we show in Table 3. Generally, on the critical side, they noted that if a student answer is too short or incomplete, neither model explicitly flags the missing details. More specifically, Qwen2.5-0.5B was praised for clarity and concision, whereas Llama-3.2-3B tended to repeat assignment details without offering actionable guidance. Annotators observed that Llama-3.2-1B often gave more specific and constructive feedback but occasionally sounded harsh, while Llama-3.1-8B sometimes overlooked key aspects. Overall, although Qwen2.5-14B achieved high win rates (94, 77, 81 across three annotators), these insights suggest that even top-performing models could improve in error detection, tone refinement, and contextual sensitivity.
To further quantify, in Figure 5, we plot for each model the net balance of optional rater comments in five qualitative categories: Actionability (AC), Goal Orientation (GO), User Friendliness (UF), Consistency (CO) and Student Autonomy (AY) (Carless et al. 2011; Wiggins 2012). Squares denote the base models and circles the SEFL-tuned variants. Horizontal whiskers give bootstrap Wilson 95% confidence intervals for the net balance. We compute these intervals on the proportion of positive remarks and then transform them to the net scale via , where is the total number of comments and the positive proportion. Annotators added comments only when a response stood out, so the plot reveals both the direction and the strength of impressions. We show several examples of annotations in Table 4. The annotators also had a choice if the feedback was non-nonsensical, which happened around 12% of the time mainly with the smaller sized models.
The two smallest models, Qwen2.5-0.5B and Llama-3.2-1B, receive more negative than positive remarks on Consistency and Goal Orientation, matching earlier findings that they sometimes drift from the student answer or overlook core requirements. Llama-3.1-8B shows the only clearly positive balance in Actionability and User Friendliness, yet its interval for Consistency still lies below zero. Qwen2.5-14B gains more favorable notes on tone and clarity than the smaller models but still shows a negative alignment gap. With respect to Student Autonomy, there are mostly neutral comments.
| Method/Dataset | J1 | J2 | J3 | J4 |
| Book2Dial | 43 | 37 | 60 | 27 |
| SEFL | 57 | 63 | 40 | 73 |
6.2 LLM-as-a-Judge
We used LLM judges to rate the feedback generated by SEFL-tuned models against their vanilla counterparts. This provides a rapid, scalable way to measure feedback quality, reducing the need for extensive human annotation. We let the LLM judges rate the same examples as the humans annotated. As shown in Figure 2, three out of four LLM judges consistently favored SEFL-tuned Qwen2.5-0.5B, Llama-3.2-1B, and Llama-3.2-3B. With Command-R, we notice that it performs worse than GPT-4o and Claude-3.5-Sonnet on JudgeArena. This indicates that the misalignment might have to do with instruction following capabilities of these models, this is highlighted by Command-R being weaker in Chatbot Arena (Zheng et al. 2024). Nonetheless, we see it as a practical first step for large-scale feedback comparisons in educational contexts. We recommend supplementing LLM-based assessments with targeted human evaluations for more granular insights, possibly aligning more with instructional objectives or possibly even fine-tuning LLMs with rubrics for better judges of long form text (Kim et al. 2024).
6.3 Comparison to Prior Work
The work that is closest to ours is Book2Dial (Wang et al. 2024a). It is a framework that turns textbooks into synthetic conversations between a student model and a teacher model. The student only sees high-level cues such as section titles or key concepts, while the teacher has full access to the source passage, prompting a question-answer exchange that stays aligned with the book content. Instead, we focus on any type of assignment not limited to textbooks.
To compare the two methods, we further fine-tune Qwen2.5-14B with the Book2Dial data in the same way as our SEFL-tuned version. We use the existing Book2Dial data consisting of 889 dialogues, we preprocess the data such that there is only a one-turn conversation, resulting in 5.3K conversation pairs. We then run the fine-tuned model over the same samples evaluated in Subsection 4.2 and evaluate them with the same 4 judges and evaluation criteria.444At the time of writing, we ran out of funding for the human annotators and thus compare here only with LLM-as-a-Judge.
Table 5 shows that three out of four judges prefer SEFL, yielding an average win rate of 58%. Only command-r-plus leans the other way with a 60% win rate for Book2Dial, but an average of 42% win rate. The finding is consistent with its earlier drift of seemingly failing to instruction follow. The results confirm that SEFL produces higher quality feedback than the textbook based dialogues in Book2Dial.
7 Conclusion
We introduced SEFL, a framework that simulates teacherstudent interactions via two-agent LLMs to generate synthetic data for fine-tuning smaller models. Empirical and qualitative results show the potential of this framework for enhancing feedback in educational settings. This approach yields concise, context-sensitive feedback that often surpasses original instruction-tuned models under both LLM-as-a-judge and human evaluations. Yet human insights remain indispensable for capturing nuances like clarity and tone. SEFL provides a promising avenue for immediate, personalized feedback at scale, extending beyond the educational domain.
Limitations.
We acknowledge that SEFL relies on synthetically generated assignments and errors, and are not real student submissions, which could have implications. Although this approach helps create large datasets, it risks producing feedback unaligned with authentic classroom contexts. Our evaluation also uses LLM-based judges, introducing potential biases related to each judge’s training data and objectives. Lastly, while we focused on short-answer tasks, longer or more domain-specific assignments may require specialized or more diverse synthetic data.
Ethical Statement
The use of synthetic data provides an opportunity to train automated feedback systems without the constraints of privacy and consent that come from repurposing actual student assignments and teacher feedback as training data. However, it also raises questions about transparency and potential misuse (Lindsay et al. 2024). For instance, malicious actors could manipulate synthetic data to disseminate misleading or biased feedback, undermining trust in educational tools. Users may also mistake synthetic feedback for real, expert guidance. Moreover, automated feedback systems risk reinforcing biases if the underlying models carry skewed training data. We believe educators and institutions should remain aware of these risks and incorporate human oversight to ensure that such systems complement, rather than replace, genuine pedagogical engagement with real teachers.